Specify the oxidation numbers of the metals in the following coordination entities: (i) [Co(H2O)(CN)(en)2]2+ (iii) [PtCl4]2– (v) [Cr(NH3)3Cl3] (ii) [CoBr2(en)2]+ (iv) K3[Fe(CN)6]

Grade:9

2 Answers

Gaurav
9 years ago
Hello Student
[Co(H2O)(CN)(en)2]2+ : x +0+(-1) + 0 = +2 or x = +1
[PtCl4]2– : x+( -4) = -2 or x = +2
[Cr(NH3)3Cl3] : x+0+-3 = 0 or x= +3
[CoBr2(en)2]+ : x+ 2(-1) +0 = +1 or x = +1
K3[Fe(CN)6 : -3 = x+ (-6) or x = +3
Anjali kumari
5 years ago
The answer is  +1, +2, +3 ,+1 ,+3 respectively. Adbhud soc six dosto udyogo yuddh even sharm yahi ic sk icn ushn us dy.